An investor owns a portfolio consisting of two mutual funds, A and B, with 50% invested in A. The following table lists the inputs for these funds.

Measures Fund A Fund B
Expected value 10    3   
Variance 63 26
Covariance 28

  
a. Calculate the expected value for the portfolio return. (Round your answer to 2 decimal places.)


b.
Calculate the standard deviation for the portfolio return. (Round intermediate calculations to at least 4 decimal places. Round your final answers to 2 decimal places.)
Answer #1

a)

expected value for the portfolio return =0.5*10+0.5*3=6.5

b)

standard deviation =sqrt(0.52*63+0.52*26+2*0.5*0.5*28)=6.02